OpenStudy (smurfy14):

Help please!! A bike accelerates uniformly from rest to a speed of 7.10 m/s over a distance of 25.4 m. Determine the acceleration of the bike.

OpenStudy (nikita2):

S = v0*t + (a*t^2)/2 and v = v0 + a*t, in your case v0 = 0 So S = (at^2)/2, v= at, -> t = v/a S = (a*(v/a)^2)/2 25.4 = ((7.1)^2)/2a ->a = ((7.1)^2)/2*25.4 = 0.992 m/s^2